# A Nonlocal Aw-Rascle-Zhang system with linear pressure term

### v1

@@ Abstract.
In this paper, we study a nonlocal extension of the Aw-Rascle-Zhang traffic model, where the pressure-like term is modeled as a convolution between vehicle density and a kernel function. This formulation captures nonlocal driver interactions and aligns structurally with the Euler-alignment system studied in [23]. Using a sticky particle approximation, we construct entropy solutions to the equation for the cumulative density and prove convergence of approximate solutions to weak solutions of the nonlocal system. The analysis includes well-posedness, stability estimates, and an entropic selection principle.
Keywords: Nonlocal Aw-Rascle-Zhang system, sticky particle dynamics, weak solutions
@@ 1. Introduction
Then, the nonlocal ARZ system (1.2) can be rewritten in the Euler-alignment form (1.7).
@@
The purpose of this paper is to prove the well-posedness and stability estimates for system (1.5), and to provide an entropic selection principle. In particular, we develop a global well-posedness theory for weak solutions of (1.5) with measure-valued density and bounded velocity. Our approach is based on an adaptation of the sticky particle approximation, originally introduced by Brenier and Grenier [5] to analyze the 1D pressureless Euler equations and developed in [23]. The novelty with respect to the previous results in [23] is that we are considering an integrable and bounded kernel function $\phi$ which is not necessarily symmetric (even) nor positive. The shape of its antiderivative $\omega$, see (1.6) and (H), can be chosen closer to that one of the anisotropic kernel functions usually used in literature for nonlocal traffic flow models such as in [9, 10, 11]. The choice of a non-symmetric kernel function makes the analysis more involved and interesting, moreover it is coherent from the modeling point of view because drivers are supposed to adapt the velocity with respect to what happens mainly downstream in the traffic. It is worth quoting the following recent paper [19] in which the authors consider a nonlocal version of the ARZ model with nonlocality in the pressure density obtained as a convolution type term between a positive, bounded and non-increasing on $0,\infty)$ kernel function. In particular, [19] is focused on the well-posedness of smooth solutions and the phase-plane analysis. Other recent related papers are [25] and [1]. In the first paper the authors establish existence, uniqueness and Oleĭnik estimates, together with the singular limit, for a nonlocal version of the so-called Generalized Aw-Rascle-Zhang model (GARZ) [15], where the drivers’ velocity depends on both the empty road velocity and the convolution of the density with an anisotropic kernel. In [1] the authors analyze the well-posedness and the singular limit for another nonlocal version of the GARZ model in which the nonlocality is present in the flux function as a convolution-type term between the velocity function and an anisotropic kernel function.
@@ 2. Sticky particle Cucker-Smale
(2.14) $\underline{\psi}-\mathinner{\!\left\lVert\omega\right\rVert}_{L^{\infty}(\mathbb{R})}\leq v_{i}(t)\leq\bar{\psi},\quad\forall i=1,\dots,N,\quad t\geq 0.$
@@
$\tilde{M}\mathrel{\mathop{\ordinarycolon}}=\max\{\,|\underline{\psi}-\mathinner{\!\left\lVert\omega\right\rVert}_{L^{\infty}(\mathbb{R})}|,\,|\bar{\psi}|\,\}.$
@@
(2.17) $\psi_{i}(t)-\sum_{j=1}^{N}m_{j}\omega(x_{i}(t)-x_{j}(t))\geq\underline{\psi}-\mathinner{\!\left\lVert\omega\right\rVert}_{L^{\infty}(\mathbb{R})}.$
@@
$\tilde{M}\mathrel{\mathop{\ordinarycolon}}=\max\{\,|\underline{\psi}-\mathinner{\!\left\lVert\omega\right\rVert}_{L^{\infty}(\mathbb{R}))}|,|\bar{\psi}|\}.$
@@
(2.18) $\frac{\sum_{j=i_{*}(t)}^{k}m_{j}\psi_{j}(t-)}{\sum_{j=i_{*}(t)}^{k}m_{j}}\geq\frac{\sum_{j\in J_{i}(t)}m_{j}\psi_{j}(t-)}{\sum_{j=\in J_{i}(t)}m_{j}}=\psi_{i}(t+)\geq\frac{\sum_{j=k}^{i^{*}(t)}m_{j}\psi_{j}(t-)}{\sum_{j=k}^{i^{*}(t)}m_{j}}.$
@@
Writing explicitly the sums, what we obtain is
@@ 3. The scalar balance law
In this section we introduce a scalar balance law connected to system (1.5). Inspired by the work of Brenier and Grenier [5] and the work of Leslie and Tan [23], we aim to prove the well-posedness of a scalar balance law linked to (1.5), and use the entropy condition of this scalar balance equation as selection principle for the solution of (1.5). To do this, we define two new quantities $M$ and $Q$, which are respectively the antiderivatives of $\rho$ and $\rho\psi$, namely
@@
Integrating both equations in (1.5), it is simple to verify that the two quantities above satisfy respectively the following equations
@@
We observe that $M$ is a monotone function, being $\rho(x,t)$ positive, then it is possible to define a map $A$, depending only on the initial data $(M^{0},Q^{0})$, such that $A(M^{0})=Q^{0}$ and if $(x,t)$ lies on a characteristic path generated by $(x_{0},0)$ with velocity field $u$, then
@@
Using this last relation and the definition of $\psi$, we get
@@
We can substitute this last equation in the transport equation of $M$, see (3.2): and we obtain our scalar balance law for the quantity $M$,
(3.3) $\partial_{t}M+\partial_{x}(A(M))=(\phi\ast M)\partial_{x}M.$
In these steps we use the fact that
$\phi\ast M=\omega\ast\rho,$
and this relation holds true, indeed
@@
Now, we want to define in which weak sense the scalar balance law must be satisfied. It is possible to define a couple entropy/entropy flux $(\eta,q)$ such that
@@
with $\eta$ Lipschitz and convex, the associated entropy inequality is
@@
where $\varphi\in C_{c}^{\infty}(\mathbb{R}\times[0,\infty))$ is a test function. The first integral of the inequality can be interpreted in the sense of distribution, while in the second term $\partial_{x}(\eta(M))$ is a Radon measure. Thanks to [6, Th. 4.19 (Young)] we know that the term $\phi\ast M$ is in $L^{\infty}(\mathbb{R})\cap C^{0}(\mathbb{R})$.
It is possible to notice that the integral
@@
is well defined, indeed $\partial_{x}(\eta(M))$ is a Radon measure and $\phi\ast M\in L^{\infty}(\mathbb{R})\cap C^{0}(\mathbb{R})$, but, by definition $\varphi\in C_{c}^{\infty}(\mathbb{R})$, and then the product $\varphi(\phi\ast M)$ has compact support.
@@
Thus, this means that the distributional derivative of a $BV$ function $u$ is an $\mathbb{R}^{N}$-vector valued Radon measure. Using this result we can consider (3.5) equals to
@@
and so we can consider (3.4) in the sense of distributions, then we give the following definition.
@@
Let $M\mathrel{\mathop{\ordinarycolon}}\mathbb{R}\times[0,T]\to\mathbb{R}$, with $T>0$, we say that $M$ is an entropy weak solution of (3.3) if it satisfies the following conditions:
  • the inequality (3.4) is satisfied for any pair $(\eta,q)$ entropy/entropy-flux in the sense discussed above;
@@
It is possible to extend Definition 3.2 of weak entropy solution for $M(x,t)$ to $\mathbb{R}\times\mathbb{R}_{+}$, requiring that for any constant $\bar{T}>0,$ $M(x,t)$ is a weak entropy solution in $\mathbb{R}\times[0,\bar{T}]$.
In particular, in the equation (3.4) we will consider Kružkov entropy pairs, see [21], namely
(3.6) $\eta(M)=|M-\alpha|,\quad q(M)=\text{sgn}(M-\alpha)(A(M)-A(\alpha)),$
where $\alpha\in\mathbb{R}$.
Now, we can state the Rankine-Hugoniot condition and the Oleinik entropy condition. Assume that $M$ takes the values $M_{l}$ and $M_{r}$ on the two sides of a shock along a curve $C=\{(x,t)\mathrel{\mathop{\ordinarycolon}}x=x(t)\}$, and denote by $s(t)=\dot{x}(t)$ the shock-speed. The entropy condition becomes
@@
where $n=(n_{1},n_{2})=\frac{1}{\sqrt{1+s^{2}}}(1,-s)$ and with the notation $f]=f(M_{l})-f(M_{r})$. Substituting $n_{1}$ and $n_{2}$ in (3.7), we get
@@
To get the Oleinik entropy condition, we consider the following. If $M_{l}<M_{r}$, we take $\eta(m)=(m-\theta)H(m-\theta)$ and $q(m)=(A(m)-A(\theta))H(m-\theta)$ with $\theta\in(M_{l},M_{r}),$ where the function $H$ is the right-continuous Heaviside function, i.e.
@@
Now we aim to prove the existence of a solution for the scalar balance law (3.3) and we start to study the following discretized formulation
@@
with initial datum
(3.12) $M_{N}(x,0)=M_{N}^{0}(x),$
where the flux function $A_{N}$ is a continuous and piecewise linear function with breakpoints at $\theta_{i}$ for $i=1,\dots,N-1$:
(3.13) $A_{N}\mathrel{\mathop{\ordinarycolon}}[0,1]\rightarrow\mathbb{R},\quad\mbox{$A_{N}^{\prime}$ is constant in each interval $(\theta_{i-1},\theta_{i}$), $\forall i\in\{1,\dots,N\}$.}$
The main purpose of this section is to prove that we can provide an entropy solution to (3.11)-(3.12) through the sticky particle Cucker-Smale dynamics.
First of all, we state some assumptions, in particular we consider a piecewise constant initial datum $M_{N}^{0}$ of the form
(3.14) $M_{N}^{0}(x)=\sum_{j=1}^{N}m_{j}H(x-x_{j}^{0}).$
We also assume that $x_{1}^{0}\leq x_{2}^{0}\leq...\leq x^{0}_{N}$ are the initial positions of particles in the sticky particle Cucker-Smale dynamics (2.1)-(2.5)-(2.6), while we assume the mass $m_{j}>0$ for any $j=1,\dots,N$ and ${\sum_{j=1}^{N}}m_{j}=1$.
@@
Under these hypotheses, it is possible to prove the existence of an entropy weak solution of the discretized scalar balance law coupled with an initial datum (3.11)-(3.12), as in the following Theorem 3.3.
@@
Consider the scalar balance law (3.11) with initial datum (3.12) of the form given in (3.14) and with a flux function $A_{N}$ satisfying (3.13). We define
@@
is an entropy solution of the discretized Cauchy problem (3.11)-(3.12) satisfying the following:
@@
    at points of discontinuity the following equality holds true
@@
    at the discontinuity points the following inequality holds true
@@
    with $\theta\in(\theta_{i_{*}(t)-1},\theta_{i^{*}(t)})$. This means that an Oleinik entropy condition is satisfied.
Thus, $M_{N}(x,t)$ is an entropy solution of (3.11).
@@
We start to prove (a). This point holds true by construction. Indeed, we know that
@@
and, by Proposition 2.2, we also know $\{x_{i}(t)\}_{i=1}^{N}\subset[-R(T),R(T)],$ then (a) holds true.
We continue with the proof of (b). The case $x<x_{1}(t)$ is trivial. For any $x\geq x_{1}(t)$, let us consider the highest index $k\in\{1,\dots,N\}$, such that $x_{k}(t)\leq x<x_{k+1}(t)$, then we can write
(3.22) $A_{N}(M_{N}(x,t))=\sum_{j=1}^{k}(A_{N}(\theta_{i})-A_{N}(\theta_{i-1}))=\sum_{j=1}^{k}m_{j}\psi_{i}^{0}.$
Moreover, from (2.8) we have
@@
Substituting (3.23) in (3.22), we get (3.19).
Concerning the Rankine-Hugoniot condition in (c) and the Oleinik entropy condition in (d), we proceed as in the following.
By construction, we can see that the function $M_{N}(\cdot,t)$ is piecewise constant with discontinuities only along the curves $C_{i}=\{(x_{i}(t),t)\,\mathrel{\mathop{\ordinarycolon}}\,t\geq 0\}$. Furthermore, we observe that $M_{N}(x,t)$ takes values in the set $\{0,\theta_{1},\ldots,\theta_{N-1},1\},$ therefore $A_{N}(M_{N}(x,t))$ is well defined outside the curves $C_{i}$. Thus, we need to verify the conditions along these discontinuities with shock speed $\sigma_{i}(t)=v_{i}(t)$. Let us fix a time $t$ and a point $(x_{i}(t),t)$ on the curve $C_{i}$. By definition of $M_{N}$ in (3.18), we obtain
@@
where the last two equalities are obtained using the relations (3.16) and by iterating (2.8) across interactions. Thanks to [6, Th. 4.19 (Young)], the function $\phi\ast M_{N}\in L^{\infty}(\mathbb{R})\cap C^{0}(\mathbb{R})$, and
@@
Putting together (3.24) and (3.25), and using (3.26), we end up with
@@
then the Rankine-Hugoniot condition in (c) is proved. Now, we check (3.21): we have
@@
with $\theta\in(\theta_{i_{*}(t)-1},\theta_{i^{*}(t)})$. Being $A_{N}$ piecewise linear, it is sufficient to prove that condition (3.27) is verified in the breakpoints, namely $\theta=\theta_{k}$ with $k\in\{\theta_{i_{*}(t)-1},\dots,\theta_{i^{*}(t)}\}$. Then, we can write
@@
where the inequality sign is due to Lemma 2.3. ∎
The next step is to study the solution of the scalar balance law (3.3) using the result obtained in Theorem 3.3 for the discretized solution $M_{N}(x,t)$. The idea is to pass to the limit for $N$ that goes to $\infty$. Thus, let us analyze the following Cauchy problem

### v2

@@ Abstract.
In this paper, we study a nonlocal extension of the Aw-Rascle-Zhang traffic model, where the pressure-like term is modeled as a convolution between vehicle density and a kernel function. This formulation captures nonlocal driver interactions and aligns structurally with the Euler-alignment system studied in [23]. Using a sticky particle approximation, we construct entropy weak solutions to the equation for the cumulative density and prove convergence of approximate solutions to weak solutions of the nonlocal system. The analysis includes well-posedness, stability estimates, and an entropic selection principle.
Key words and phrases:
Nonlocal Aw-Rascle-Zhang system, sticky particle dynamics, weak solutions
@@ 1. Introduction
Then, the nonlocal ARZ system (1.2) with linear pressure can be rewritten in the Euler-alignment form (1.7).
@@
The purpose of this paper is to prove the well-posedness and stability estimates for system (1.5), and to provide an entropic selection principle. In particular, we develop a global well-posedness theory for weak solutions of (1.5) with measure-valued density and bounded velocity. Our approach is based on an adaptation of the sticky particle approximation, originally introduced by Brenier and Grenier [5] to analyze the 1D pressureless Euler equations and developed in [23]. The novelty with respect to the previous results in [23] is that we are considering an integrable and bounded kernel function $\phi$ which is not necessarily symmetric (even) nor positive. The shape of its antiderivative $\omega$, see (1.6) and (H), can be chosen closer to that one of the anisotropic kernel functions usually used in literature for nonlocal traffic flow models such as in [9, 10, 11].
The choice of a non-symmetric kernel function makes the analysis more involved and interesting; moreover it is coherent from the modeling point of view because drivers are supposed to adapt the velocity with respect to what happens mainly downstream in the traffic. It is worth quoting the following recent paper [19] in which the authors consider a nonlocal version of the ARZ model with nonlocality in the pressure density obtained as a convolution type term between a positive, bounded and non-increasing on $0,\infty)$ kernel function. In particular, [19] is focused on the well-posedness of smooth solutions and the phase-plane analysis.
Other recent related papers are [25] and [1]. In the first paper the authors establish existence, uniqueness and Oleĭnik estimates, together with the singular limit, for a nonlocal version of the so-called Generalized Aw-Rascle-Zhang model (GARZ) [15], where the drivers’ velocity depends on both the empty road velocity and the convolution of the density with an anisotropic kernel. In [1] the authors analyze the well-posedness and the singular limit for another nonlocal version of the GARZ model in which the nonlocality is present in the flux function as a convolution-type term between the velocity function and an anisotropic kernel function.
@@ 2. Sticky particle Cucker-Smale
(2.15) $\underline{\psi}-\|\omega\|_{L^{\infty}(\mathbb{R})}\leq v_{i}(t)\leq\bar{\psi},\quad\forall i=1,\dots,N,\quad t\geq 0.$
@@
$\tilde{M}\mathrel{\mathop{\ordinarycolon}}=\max\{\,|\underline{\psi}-\|\omega\|_{L^{\infty}(\mathbb{R})}|,\,|\bar{\psi}|\,\}.$
@@
(2.18) $\psi_{i}(t)-\sum_{j=1}^{N}m_{j}\omega(x_{i}(t)-x_{j}(t))\geq\underline{\psi}-\|\omega\|_{L^{\infty}(\mathbb{R})}.$
@@
$\tilde{M}\mathrel{\mathop{\ordinarycolon}}=\max\{\,|\underline{\psi}-\|\omega\|_{L^{\infty}(\mathbb{R})}|,|\bar{\psi}|\}.$
@@
(2.19) $\frac{\sum_{j=i_{*}(t)}^{k}m_{j}\psi_{j}(t-)}{\sum_{j=i_{*}(t)}^{k}m_{j}}\geq\frac{\sum_{j\in J_{i}(t)}m_{j}\psi_{j}(t-)}{\sum_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}j\in J_{i}(t)}}m_{j}}=\psi_{i}(t+)\geq\frac{\sum_{j=k}^{i^{*}(t)}m_{j}\psi_{j}(t-)}{\sum_{j=k}^{i^{*}(t)}m_{j}}$
where $i_{*}(t)$, $i^{*}(t)$ are defined in (2.5).
@@
Writing explicitly the sums, we have
@@ 3. The scalar balance law
Inspired by the work of Brenier and Grenier [5] and the work of Leslie and Tan [23], we prove the well-posedness of a scalar balance law linked to (1.5), and we use its entropy condition as selection principle for the uniqueness of the solution to (1.5). To do this, we define two new quantities, which are respectively the antiderivatives of $\rho$ and $\rho\psi$,
@@
Integrating (1.5) in the space variable, we verify that $M$ and $Q$ satisfy
@@
We observe that $M$ is a monotone function, being $\rho(x,t)$ non negative, and so choosing an initial data
(3.3) $(M,Q)(x,0)=(M^{0},Q^{0}),$
such that $Q^{0}$ is constant in the intervals where $M^{0}$ is also constant, it is possible to define a map $A$ such that such that $A(M^{0}(x))=Q^{0}(x)$ for any $x\in\mathbb{R}$, moreover if $(x,t)$ lies on a characteristic path generated by $(x_{0},0)$ with velocity field $u$, then
@@
Using this relation and the definition of $\psi$, we get
@@
Substituting this last equation in (3.2)1, we obtain a scalar balance law for $M$,
(3.4) $\partial_{t}M+\partial_{x}(A(M))=(\phi\ast M)\partial_{x}M,$
where we use $\phi\ast M=\omega\ast\rho.$ Indeed we have
@@
We define an entropy/entropy flux pair $(\eta,q)$ as a pair of functions $\eta=\eta(M)$ and $q=q(M)$ such that
@@
with $\eta$ Lipschitz and convex. The entropy inequality is
@@
where $\varphi\in C_{c}^{\infty}(\mathbb{R}\times[0,\infty))$ is a test function.
Remark 3.1.
The first integral of the inequality can be interpreted in the sense of distribution, while in the second term $\partial_{x}(\eta(M))$ is a Radon measure. Thanks to [6, Th. 4.19 (Young)] we know that the term $\phi\ast M$ is in $L^{\infty}(\mathbb{R})\cap C^{0}(\mathbb{R})$, then the integral
@@
is well defined by the fact that $\partial_{x}(\eta(M))$ is a Radon measure, $\phi\ast M\in L^{\infty}(\mathbb{R})\cap C^{0}(\mathbb{R})$ and $\varphi\in C_{c}^{\infty}(\mathbb{R})$, thus the product $\varphi(\phi\ast M)$ has compact support.
@@
Thanks to this result we rewrite (3.6) as
@@
and so we can consider (3.5) in distributional sense. We provide the definition of entropy weak solution to (3.3)-(3.4).
@@
Let $M\mathrel{\mathop{\ordinarycolon}}\mathbb{R}\times[0,T]\to\mathbb{R}$, with $T>0$, we say that $M$ is an entropy weak solution of (3.4)-(3.3) if it satisfies the following conditions:
  • $(M,Q)(x,0)=(M^{0},Q^{0})$;
  • the inequality (3.5) is satisfied for any entropy/entropy-flux pair $(\eta,q)$ in the sense discussed above;
@@
It is possible to extend Definition 3.3 of entropy weak solution to $\mathbb{R}\times\mathbb{R}_{+}$, requiring that for any constant $\bar{T}>0,$ $M(x,t)$ is an entropy weak solution in $\mathbb{R}\times[0,\bar{T}]$.
Now, we state the Rankine-Hugoniot condition and the Oleinik entropy condition. Assume that $M$ takes the values $M_{l}$ and $M_{r}$ on the two sides of a shock along a curve $C=\{(x,t)\mathrel{\mathop{\ordinarycolon}}x=x(t)\}$, and denote by $s(t)=\dot{x}(t)$ the shock-speed. The entropy condition becomes
@@
where $n=(n_{1},n_{2})=\frac{1}{\sqrt{1+s^{2}}}(1,-s)$, with the notation $f]=f(M_{l})-f(M_{r})$. Substituting $n_{1}$ and $n_{2}$ in (3.7), we get
@@
To get the Oleinik entropy condition, we consider the following. If $M_{l}<M_{r}$, we take $\eta(m)=(m-\theta)H(m-\theta)$ and $q(m)=(A(m)-A(\theta))H(m-\theta)$ with $\theta\in(M_{l},M_{r}),$ where the function $H$ is the right-continuous Heaviside function
@@
3.1. Construction of approximated solutions
In this subsection we investigate the existence of entropy weak solutions to the scalar balance law (3.4) and we start by studying the following discretized formulation:
@@ 3.1. Construction of approximated solutions
together with initial datum
(3.12) $M_{N}(x,0)=M_{N}^{0}(x)\,.$
Here the flux function $A_{N}$ is a continuous and piecewise linear function with breakpoints at $\theta_{i}$ for $i=1,\dots,N-1$:
(3.13) $A_{N}\mathrel{\mathop{\ordinarycolon}}[0,1]\rightarrow\mathbb{R},\quad\mbox{$A_{N}^{\prime}$ is constant in each interval $(\theta_{i-1},\theta_{i}$)},\quad i=1,\dots,N.$
The main purpose of this subsection is to construct an entropy weak solution to (3.11)-(3.12) by means of the sticky particle Cucker-Smale dynamics.
To start, let’s consider a piecewise constant initial datum $M_{N}^{0}$ as follows:
(3.14) $M_{N}^{0}(x)=\sum_{j=1}^{N}m_{j}H(x-x_{j}^{0}),$
where $x_{1}^{0}\leq x_{2}^{0}\leq...\leq x^{0}_{N}$ are the initial positions of particles in the sticky particle Cucker-Smale dynamics (2.1)-(2.6)-(2.7), and the masses are such that $m_{j}>0$ for any $j=1,\dots,N$ and ${\sum_{j=1}^{N}}m_{j}=1$.
@@
Under these hypotheses, we prove the existence of an entropy weak solution to (3.11)-(3.12) with initial data as in (3.14).
@@
Let $A_{N}$ be a flux function satisfying (3.13), and consider $M_{N}^{0}$ as in (3.14). Define
@@
is an entropy weak solution of the discretized Cauchy problem (3.11)-(3.12) satisfying the following:
@@
    at points of discontinuity $x_{i}(t)$, $i=1,\dots,N$, we have
@@
    at the points of discontinuity the following inequality holds true
@@
    with $\theta\in(\theta_{i_{*}(t)-1},\theta_{i^{*}(t)})$. Thus the Oleinik entropy condition is satisfied.
@@
We show that the function $M_{N}(x,t)$ defined in (3.18) is an entropy weak solution to (3.11)-(3.12), namely, we prove that it satisfies points (a), (b), (c), and (d) of the statement. The proof is divided into steps for a better readability.
  • (a)
    This point holds true by construction, indeed
@@
    and, by Proposition 2.2, we have $\{x_{i}(t)\}_{i=1}^{N}\subset[-R(T),R(T)],$ then (a) holds true.
  • (b)
    The case $x<x_{1}(t)$ is trivial. For any $x\geq x_{1}(t)$, let us consider the highest index $k\in\{1,\dots,N\}$, such that $x_{k}(t)\leq x<x_{k+1}(t)$, with $x_{N+1}(t)=+\infty$, then we have
    (3.22) $A_{N}(M_{N}(x,t))=\sum_{j=1}^{k}(A_{N}(\theta_{i})-A_{N}(\theta_{i-1}))=\sum_{j=1}^{k}m_{j}\psi_{i}^{0},$
    moreover, from (2.9) we get
@@
    Substituting (3.23) in (3.22), we obtain (3.19).
  • (c)
    We can see by construction that the function $M_{N}(\cdot,t)$ is piecewise constant with discontinuities only along the curves $C_{i}=\{(x_{i}(t),t)\,\mathrel{\mathop{\ordinarycolon}}\,t\geq 0\}$. Furthermore, we observe that $M_{N}(x,t)$ takes values in the set $\{0,\theta_{1},\ldots,\theta_{N-1},1\},$ therefore $A_{N}(M_{N}(x,t))$ is well defined outside the curves $C_{i}$. Thus, we need to verify the conditions along these discontinuities with shock speed $\sigma_{i}(t)=v_{i}(t)$. Let us fix a time $t$ and a point $(x_{i}(t),t)$ on the curve $C_{i}$. By definition of $M_{N}$ in (3.18), we obtain
@@
    where the last two equalities are obtained using the relations (3.16) and by iterating (2.9) across interactions. Thanks to [6, Th. 4.19 (Young)], the function $\phi\ast M_{N}\in L^{\infty}(\mathbb{R})\cap C^{0}(\mathbb{R})$ and can be expressed as follows:
@@
    Putting together (3.24), (3.25), and using (3.26), we obtain
@@
    then the Rankine-Hugoniot condition is proved.
  • (d)
    Now, we check (3.21): we have
@@
    with $\theta\in(\theta_{i_{*}(t)-1},\theta_{i^{*}(t)})$. Being $A_{N}$ piecewise linear, it is sufficient to prove that (3.27) is verified at the breakpoints $\theta=\theta_{k}$ with $k\in\{i_{*}(t),\dots,i^{*}(t)-1\}$. Then, we can write
@@
    where the inequality is due to Lemma 2.3.
3.2. The solution of the scalar balance law
The next step is to study the solution of the scalar balance law (3.4) using the result obtained in Theorem 3.4. The idea is to pass to the limit as $N$ goes to $\infty$. Let us consider the Cauchy problem